PROBLEM:

On January 1, 2019 Day Co. leased a new machine from Parr with the following pertinent information:

Lease term                                                                                       6 years

Annual rental payable at the beginning of each year        $50,000

Useful life of machine                                                                  8 years

Day’s incremental borrowing rate                                           15%

Implicit interest rate in lease (known by Day)                     12%

Present value of annuity of 1 in advance for 6 periods at

            12%                                                                                        4.61

            15%                                                                                        4.35

The lease passes ownership of the machine to Day at the termination of the lease.  The cost of the machine on Parr’s accounting record is $375,500.  

QUESTIONS:

  1. At the beginning of the lease term, Day should record a lease liability of
  1. $375,500
  2. $230,500
  3. $217,500
  4. $0
  1. Is this a capital or operating lease?
  1. Record/Journalize the entry Day should make at the inception/beginning of the lease Jan 1, 2019
  1. Record/journalize Day’s lease payment at 1/1/ 2019

  1. Record/journalize Day’s annual interest payment  for 2019 (use the effective interest rate method)

  1. How much is Day’s lease liability (how much he owes) at the end of 2019.
  1. Record/journalize the depreciation @ the end of 2019.
Answer #1
Given Information
Useful life of Machine 8 years
Lease Term 6 years
Annual Rent for 6 years $   50,000.00
Implicit Interest Rate for lessee 12%
Present Value Annuity factor @ 12% 4.61
Present Value of all the future Lease Rent Payments
= Annual Lease Rental * PF Factor
= 50000 * 4.61 = 230500
So, at the beginning of the lease term, the lease expense should be recorded at $ 2,30,500.00
To qualify as a capital lease, a lease contract must satisfy any of the four criteria.
First, the life of the lease must be 75% or greater for the asset's useful life.
Second, the lease must contain a bargain purchase option for a price less than the market value of an asset.
Third, the lessee must gain ownership at the end of the lease period.
Finally, the present value of lease payments must be greater than 90% of the asset's market value.
In the given case the life of lease is 6 years and the useful life of the given Asset is 8 years. Which means that the Lease life is 75% of the Asset Life. Hence, this is a type of Operating Lease.
Date Journal Entries in the books of Day Co. Dr Amount Cr Amount
01.01.2019 Leased New Machine A/c Dr. $ 230,500.00
       To Lease Liability A/c $ 230,500.00
( Being Lease Liability Created )
01.01.2019 Lease Expense A/c Dr. $   50,000.00
      To Bank A/c $    50,000.00
( Being lease installment paid )
31.12.2019 Lease Liability A/c Dr. $   50,000.00
      To Right to Use Asset A/c $    50,000.00
( Being lease liability Reduced )
The Balance of Lease Liability = 230500 - 50000 = $ 180,500.00
31.12.2019 Depreciation Expense on Lease Machine A/c Dr. $   38,417.00
      To Accumulated Depreciation on Leased Machine $    38,417.00
( Being depreciation on Leased Machine charges as per SLM
= 230500 / 6   = 38417 per year )
Year Op Bal Installment Amt Principal Interest Closing
1 $ 230,500.00 $    50,000.00 50000 0 $ 180,500.00
2 $ 180,500.00 $    50,000.00 $ 28,340.00 $ 21,660.00 $ 152,160.00
3 $ 152,160.00 $    50,000.00 $ 31,740.80 $ 18,259.20 $ 120,419.20
4 $ 120,419.20 $    50,000.00 $ 35,549.70 $ 14,450.30 $    84,869.50
5 $   84,869.50 $    50,000.00 $ 39,815.66 $ 10,184.34 $    45,053.84
6 $   45,053.84 $    50,000.00 $ 44,593.54 $    5,406.46 $          460.31
6 $         460.31 $        460.31 $          55.24 $                   -